OpenStudy (anonymous):

you can put a fence around a regular lot. The length of the lot must be at least 60ft. The cost of the fence along the length is $1.50 per foot. And the width is 2$ per foot. The total cost can not excede $360.

OpenStudy (anonymous):

use two variables to write a system of inequalities that models the problem what is the maximum width if the length is 60 feet?

OpenStudy (anonymous):

@Hero alittle help?

OpenStudy (anonymous):

ok, so x is the length. y is the width. x = 60ft (60*1.5)+(y*2) <= 360

OpenStudy (anonymous):

ok thanx you know how the first one is done?

OpenStudy (tkhunny):

\(x \ge 60\;ft\) \((2x(1.5))+((2y)2)) \le360\) Also, this is a "regular" lot. This might mean \(x = y\). It's not as clear as I'd like it to be.

OpenStudy (anonymous):

ok there is one more so instead of the length having a limit what if the width had a limit of 36 ft? same kind you just did tomo

OpenStudy (tkhunny):

If x = y, it is the same sort of problem. Notice, however that in tomo's setup, there is only one length and one width. In mine, there are two of each. In my opinion, the unclear problem statement makes it difficult to tell which is the correct interpretation.

OpenStudy (anonymous):

you can put a fence around a regular lot. The length of the lot must be at least 60ft. The cost of the fence along the length is $1.50 per foot. And the width is 2$ per foot. The total cost can not excede $360 what is the maximum length if the width is 36ft?

OpenStudy (tkhunny):

I get it. Still just as unclear. Is a "regular" lot "square"? Is there a side or two missing? Perhaps there was a drawing?

OpenStudy (anonymous):

no drawing itz rectangle

OpenStudy (tkhunny):

Well, there you go. \(y \le 36\;ft\)

OpenStudy (anonymous):

you don't need to know whether it's a square or a rectangle the maximum length could make a square, it could maybe make a rectangle. All it asks is to maximize the length given the restriction and an equation. Tomo put it correctly. The max length would be obtained by using the most money therefore you can make it an equality. The restriction is simply a guideline it seems